OpenStudy (anonymous):

Help please: Use the distance formula to show that the triangle ABC, with vertices A(-1,-2),B(3,2), and C(1,4), is a right triangle. Explain your reasoning.

OpenStudy (anonymous):

AB(4; 4) BC(-2; 2) AC(2; 6) Length of AB = sqrt(4^2 + 4^2) = sqrt(32) = 4sqrt(2) Length of BC = sqrt(-2^2 + 2^2) = sqrt(8) = 2sqrt(2) Length of AC = sqrt(2^2 + 6^2) = sqrt(40) = 2sqrt(10) According to pythagoras a^2 + b^2 = c^2 if it's a right triangle. 4sqrt(2) + 2sqrt(2) = 2sqrt(10) 32 + 8 = 40 And that is true. Win.

OpenStudy (anonymous):

Thank you. So basically, you find the distance between AB,BC, and AC. Then you plug it in the a2+b2=c2 formula?

OpenStudy (anonymous):

Since you can only use the distances I can't really see another way. So yeah, 'a^2 + b^2 = c^2 must be true for right triangles' is basically all you need to explain imo : )

OpenStudy (anonymous):

okay, thanks.